OpenStudy (anonymous):

Give a formula for (ABx)^T, where x is a vector and A and B are matrices of appropriate sizes.

OpenStudy (anonymous):

Do you understand the question better?

OpenStudy (anonymous):

I don't know if they want you to find an example or come up with a general solution.

OpenStudy (anonymous):

I think that it is asking for a general example because it sounds like there are multiple ways to do this. But then again I'm not sure

OpenStudy (anonymous):

Why not just let A and B be identity matrices and let x be some vector?

OpenStudy (anonymous):

the only other thing I can think to do is... \[ (ABx)^T=b\\ ABx = b^T\\ x = (AB)^{-1}b^T\\ x = B^{-1}A^{-1}b^T \]

OpenStudy (anonymous):

Does that go back to the identity matrices? or is that talking about inverses?

OpenStudy (anonymous):

The \(^{-1}\) means inverse.

OpenStudy (anonymous):

thanks!